# Floquet Nonadiabatic Mixed Quantum-Classical Dynamics in Laser-Dressed Solid Systems

### v1

@@ III.1 Electronic populations
In this analysis, we consider a lattice composed of $N=10$ sites while adhering to the parameters $J_{0}=0.3$, $\hbar\omega=0.3$, $g=0.5$, and $T=0.5$ in the context of the Holstein model and the Peierls model. To obtain these numerical results, we employ a 4th-order Runge-Kutta algorithm to propagate both classical and quantum coordinates. The chosen time step for the algorithm is $dt=0.01$. For the sake of statistical accuracy, the outcomes have been averaged across $10,000$ initial thermal conditions that conform to the Boltzmann distribution for the classical coordinates. Besides, we set the initial probability that the electron appears in each electronic state to be exactly equal in real space; thus, the probability is completely gathered in the $k=0$ state in reciprocal space. Here we graph $P_{0}(t)\equiv\left|\langle 0|\Psi(t)\rangle\right|^{2}$ as a function of time.
@@
Figure 2: Electronic population dynamics at $k=0$ under varying periodic drivings for the Holstein model, employing four methods: FMD and FSH in both real- and reciprocal-space. In scenario (a), we fix the driving frequency $\Omega=0.5$ and vary the driving amplitude $J_{1}$ from 00 to $1$. In scenario (b), with a constant driving amplitude $J_{1}=0.4$, the driving frequency $\Omega$ is altered from $0.2$ to $2$. The parameters are set as $J_{0}=0.3$, $\omega=0.3$, $g=0.5$, $T=0.5$, and $N=10$.
@@
In Fig. 2 (a), we fix the driving frequency $\Omega=0.5$. We observe that with an increase in the driving amplitude $A$, the steady-states of the electronic population at momentum zero initially decrease, followed by a subsequent increase. This observation suggests that periodic drivings have the capacity to transfer energy to electrons within solid-state materials, thereby rendering the electrons more dynamically active. In Fig. 2 (b), with a fixed driving amplitude $A=0.4$, we observe an intriguing pattern as the driving frequency ($\Omega$) is varied. Initially, the steady-states of $\rho_{0}$ decrease, followed by a subsequent increase until reaching a point where they converge to the same outcome as in the absence of drivings ($J_{1}=0$). Under fast drivings, where $\Omega$ is sufficiently large, the system may fail to respond to such rapid drivings. Consequently, the properties of solids revert to their original conditions under fast drivings as in the absence of periodic drivings. Next, we turn to the Peierls model to simulate electronic population dynamics under various periodic drivings, as shown in Fig. 3. Remarkably, we observe a consistent trend reminiscent of the findings in the Holstein model.
@@
Figure 3: Electronic population dynamics at $k=0$ under varying periodic drivings for the Peierls model, employing four methods: FMD and FSH in both real- and reciprocal-space. In scenario (a), we fix the driving frequency $\Omega=0.5$ and vary the driving amplitude $J_{1}$ from 00 to $1$. In scenario (b), with a constant driving amplitude $J_{1}=0.4$, the driving frequency $\Omega$ is altered from $0.2$ to $2$. The parameters are set as $J_{0}=0.3$, $\omega=0.3$, $g=0.5$, and $T=0.5$, and $N=10$.
@@ III.2 Charge mobility simulations
By employing our FMD and FSH methods, we also simulate charge mobility in a short time under different periodic drivings in real space, as shown in Figs. 4 and 5. Short-time charge mobility is essential for semiconductor devices or high-frequency electronics. For example, gallium nitride (GaN) high electron-mobility transistors have emerged as excellent power devicesMeneghesso et al. (2008); Meneghini et al. (2021). Here, we simulate the dynamics of the mean square displacement $\langle r^{2}\rangle$ of the charge in real space, which is proportional to the charge mobilityWang et al. (2010). Different from last subsection, here we initialize the electronic population at the middle site in real space and take an average of $10,000$ trajectories for each dynamic result.
For the Holstein model, we find that by adding Floquet drivings, short-time charge mobility can be enhanced with increasing the strength of the driving (see Fig. 4 (a)). When the driving frequency becomes sufficiently high, the system may fail to respond effectively to the rapid driving, resulting in charge mobility comparable to the condition without periodic drivings. In the Peierls model, we observe a similar trend to that in the Holstein model, as shown in Fig. 5.
Figure 4: Mean square displacement in real space ($\langle r^{2}\rangle$) as a function of time under different periodic drivings for Holstein model. a) We fix the driving frequency $\Omega=0.5$, and change the driving amplitude $J_{1}$ from 00 to $0.8$. We see $\langle r^{2}\rangle$ increases with enhancing the driving amplitude $A$. b) We fix the driving amplitude $J_{1}=0.4$, and change driving frequencies $\Omega$ from $0.2$ to $2$. We see $\langle r^{2}\rangle$ increases first and then decreases to a point where there is no driving. Here, $J_{0}=0.3$, $\omega=0.3$, $g^{2}=0.25$, $T=0.5$, and $N=10$.
Figure 5: Mean square displacement in real space ($\langle r^{2}\rangle$) as a function of time under different periodic drivings for Holstein model. a) We fix the driving frequency $\Omega=0.5$, and change the driving amplitude $J_{1}$ from 00 to $0.8$. We see $\langle r^{2}\rangle$ increases with enhancing the driving amplitude $A$. b) We fix the driving amplitude $J_{1}=0.4$, and change the driving frequencies $\Omega$ from $0.2$ to $2$. We see $\langle r^{2}\rangle$ increases first and then decreases to a point where there is no driving. Here, $J_{0}=0.3$, $\omega=0.3$, $g^{2}=0.25$, $T=0.5$, and $N=10$.
@@ III.3 Truncation in Brillouin zone
The advantage of using a real-space approach lies in its ability to precisely capture local interactions within a specific spatial region while keeping computational demands in check by limiting the scope of local interactions considered. Similarly, for phenomena restricted to a certain range within the Brillouin zone, a comparable strategy can be employed in reciprocal space. In scenarios where electronic carriers are localized due to either the short time scales not permitting full Brillouin zone traversal or their stabilization at band minima, such confinement can be effectively modeled. This concept has been effectively employed in the theoretical analysis of exciton and trion states in monolayer transition-metal dichalcogenides. In these studies, the analysis was focused around the K points in the Brillouin zone by implementing a specific truncation radius, thereby streamlining the computations.
@@
In FMD and FSH dynamics, a reciprocal space formulation allows for effective truncation when the initial population is concentrated at or near $k=0$. Here we set $N=20$, $J_{0}=0.6$, $\hbar\omega=0.2$, $J_{1}=0.3$, $\Omega=0.5$, $g^{2}=5$, and $T=1$. We symmetrically truncate the corner of the Brillouin zone and keep $4/5$, $3/5$, and $2/5$ sizes of the original one, and the corresponding size numbers $N$ are $16$, $12$, and $8$, respectively. Figs. 6 and 7 illustrate the temporal evolution of the whole population (Figs. 6 (a) and 7 (a)) and $P_{0}$ (Figs. 6 (b) and 7 (b)) for the Holstein model under different truncation scales, as calculated by FMD and FSH respectively. Remarkably, these methods yield accurate electronic population profiles even with significant truncation, such as at $k_{0}=\frac{2}{5}\pi$, closely matching the untruncated results, as can be seen in the Figs. 6 (c) and 7 (c).
@@
Figure 6: Truncated Electronic Population Dynamics $P_{k}(t)$ for the Holstein Model Using the FMD Method. a) Comparison of the truncated overall population dynamics over time against the untruncated results. b) Time evolution of the truncated $P_{0}(t)$ at the center of the Brillouin zone, juxtaposed with the untruncated counterpart. c) Analysis of the error $\Delta P_{0}(t)$, quantifying the deviation from the full Brillouin zone results, across various truncation radii $k_{0}$. Parameters: $J=0.6$, $\omega=0.2$, $J_{1}=0.3$, $\Omega=0.5$, $g^{2}=5$, $T=1$, and $N=20$.
@@
Figure 7: Truncated Electronic Population Dynamics $P_{k}(t)$ for the Holstein Model Using the FSH Method. a) Comparison of the truncated overall population dynamics over time against the untruncated results. b) Time evolution of the truncated $P_{0}(t)$ at the center of the Brillouin zone, juxtaposed with the untruncated counterpart. c) Analysis of the error $\Delta P_{0}(t)$, quantifying the deviation from the full Brillouin zone results, across various truncation radii $k_{0}$. Parameters: $J_{0}=0.6$, $\omega=0.2$, $J_{1}=0.3$, $\Omega=0.5$, $g^{2}=5$, $T=1$, and $N=20$.
@@
However, as observed in Fig. 8, the accuracy of large truncations ($k_{0}=\frac{2}{5}\pi$) in the FMD method for the Peierls model is less satisfactory. In contrast, as shown in Fig. 9, the FSH method demonstrates superior performance under similar truncation conditions, highlighting its advantages.
Importantly, truncation calculations in reciprocal space offer significant computational benefits. By reducing the size of the computational domain, they enable more efficient use of computational resources while still maintaining a high degree of accuracy. This approach is particularly beneficial when dealing with large systems, especially for the Floquet Hamiltonian, making it an attractive strategy for efficiently exploring electronic dynamics in complex models.
@@
Figure 8: Truncated Electronic Population Dynamics $P_{k}(t)$ for the Peierls Model Using the FMD Method. a) Comparison of the truncated overall population dynamics over time against the untruncated results. b) Time evolution of the truncated $P_{0}(t)$ at the center of the Brillouin zone, juxtaposed with the untruncated counterpart. c) Analysis of the error $\Delta P_{0}(t)$, quantifying the deviation from the full Brillouin zone results, across various truncation radii $k_{0}$. Parameters: $J=0.6$, $\omega=0.2$, $J_{1}=0.3$, $\Omega=0.5$, $g^{2}=5$, $T=1$, and $N=20$.
@@
Figure 9: Truncated Electronic Population Dynamics $P_{k}(t)$ for the Peierls Model Using the FSH Method. a) Comparison of the truncated overall population dynamics over time against the untruncated results. b) Time evolution of the truncated $P_{0}(t)$ at the center of the Brillouin zone, juxtaposed with the untruncated counterpart. c) Analysis of the error $\Delta P_{0}(t)$, quantifying the deviation from the full Brillouin zone results, across various truncation radii $k_{0}$. Parameters: $J=0.6$, $\omega=0.2$, $J_{1}=0.3$, $\Omega=0.5$, $g^{2}=5$, $T=1$, and $N=20$.
@@ IV Conclusions
Our study successfully constructs Floquet Hamiltonian for the Holstein and Peierls models, showing the huge impact of Floquet driving on Floquet replicas. The derived FMD and FSH, explored in both real and reciprocal space, yield consistent results through numerical simulations. This agreement validates the reliability of our theoretical framework, emphasizing the equivalence of real and reciprocal space outcomes.

### v2

@@ III.1 Electronic populations
In this analysis, we consider a lattice composed of $N=10$ sites while adhering to the parameters $J_{0}=0.3$, $\hbar\omega=0.3$, $g=0.5$, and $T=0.5$ in the context of the Holstein model and the Peierls model. To obtain these numerical results, we employ a 4th-order Runge-Kutta algorithm to propagate both classical and quantum coordinates. The chosen time step for the algorithm is $dt=0.01$. In this paper, parameters are unitless. When considering room temperature ($293.150.5T=$ 293.15 $K$) as a reference, the energy unit corresponds to $50.5$ $meV$, the time unit corresponds to $82.7$ $fs$. In addition, we have set the distance between the nearest cites is 5 Å. For the sake of statistical accuracy, the outcomes have been averaged across $10,000$ initial thermal conditions that conform to the Boltzmann distribution for the classical coordinates:
$P({r_{n},p_{n}})\propto\prod_{n}\exp\left(-\frac{1}{2T}(p_{n}^{2}+\omega^{2}r_{n}^{2})\right).$ (33)
Besides, we set the initial probability that the electron appears in each electronic state to be exactly equal in real space; thus, the probability is completely gathered in the $k=0$ state in reciprocal space. Here we graph $P_{0}(t)\equiv\left|\langle 0|\Psi(t)\rangle\right|^{2}$ as a function of time. Without light-matter interactions, for mix quantum-classical dynamics in solid, the original Tully’s SH works reasonably well. This finding has been proven in literature. For instance, in Ref. [23], their numerical results show that even without decoherence correction, SH matches well with numerically exact HEOM. With light-matter interactions, the Floquet theory is exact for periodic driving. Such that FSH should work well for laser-dressed solid in general that can be considered as benchmark.
@@
Figure 2: Electronic population dynamics at $k=0$ under varying periodic drivings for the Holstein model, employing four methods: FMD and FSH in both real- and reciprocal-space. In scenario (a), we fix the driving frequency $\Omega=0.5$ and vary the driving amplitude $J_{1}$ from 00 to $1$. In scenario (b), with a constant driving amplitude $J_{1}=0.4$, the driving frequency $\Omega$ is altered from $0.2$ to $2$. The parameters are set as $J_{0}=0.3$, $\omega=0.3$, $g=0.5$, $T=0.5$, and $N=10$. Here, the time unit corresponds to $82.7$ $fs$.
@@
In Fig. 2 (a), we fix the driving frequency $\Omega=0.5$. We observe that with an increase in the driving amplitude $A$, the steady-states of the electronic population at momentum zero initially decrease, followed by a subsequent increase. This observation suggests that periodic drivings have the capacity to transfer energy to electrons within solid-state materials, thereby rendering the electrons more dynamically active. In Fig. 2 (b), with a fixed driving amplitude $A=0.4$, we observe an intriguing pattern as the driving frequency ($\Omega$) is varied. Initially, the steady-states of $\rho_{0}$ decrease, followed by a subsequent increase until reaching a point where they converge to the same outcome as in the absence of drivings ($J_{1}=0$). Under fast drivings, where $\Omega$ is sufficiently large, the system may fail to respond to such rapid drivings. Consequently, the properties of solids revert to their original conditions under fast drivings as in the absence of periodic drivings. Next, we turn to the Peierls model to simulate electronic population dynamics under various periodic drivings, as shown in Fig. 3. Remarkably, we observe a consistent trend reminiscent of the findings in the Holstein model. For both Holstein and Peierls models, in the absence of light ($J_{1}=0$), meanfield dynamics (MD) can exhibit incorrect long-time population due to lack of detailed balance. That being said, it is well-known that MD can capture short-time dynamics correctly. On the contrary, SH can predict both short-time and long-time dynamics correctly. With the presence of light-matter interactions, we find that FMD can capture the long-time population well enough in certain regime. Still, one can use improved version of the MD method to deal with dynamics in solid state Zheng and Wang (2023).
@@
Figure 3: Electronic population dynamics at $k=0$ under varying periodic drivings for the Peierls model, employing four methods: FMD and FSH in both real- and reciprocal-space. In scenario (a), we fix the driving frequency $\Omega=0.5$ and vary the driving amplitude $J_{1}$ from 00 to $1$. In scenario (b), with a constant driving amplitude $J_{1}=0.4$, the driving frequency $\Omega$ is altered from $0.2$ to $2$. The parameters are set as $J_{0}=0.3$, $\omega=0.3$, $g=0.5$, and $T=0.5$, and $N=10$. Here, the time unit corresponds to $82.7$ $fs$.
@@ III.2 Charge mobility simulations
By employing our FMD and FSH methods, we also simulate charge mobility in a short time under different periodic drivings in real space, as shown in Figs. 4 and 5. Short-time charge mobility is essential for semiconductor devices or high-frequency electronics. For example, gallium nitride (GaN) high electron-mobility transistors have emerged as excellent power devicesMeneghesso et al. (2008); Meneghini et al. (2021). Here, we simulate the dynamics of the mean square displacement $\langle r^{2}\rangle$ of the charge in real space, which is proportional to the charge mobilityWang et al. (2010). Different from III.1, here we initialize the electronic population at the middle site in real space and take an average of $10,000$ trajectories for each dynamic result.
For the Holstein model, we find that by adding Floquet drivings, short-time charge mobility can be enhanced with increasing the strength of the driving (see Fig. 4 (a)). When the driving frequency becomes sufficiently high, the system may fail to respond effectively to the rapid driving, resulting in charge mobility comparable to the condition without periodic drivings. In the Peierls model, we observe a similar trend to that in the Holstein model, as shown in Fig. 5. These numerical results illustrate two key points: firstly, we can induce charge migration in solid materials by progressively increasing the light intensity; secondly, when the light intensity is fixed, it is essential to identify the optimal intensity (neither too high nor too low) to maximize charge mobility.
Figure 4: Mean square displacement in real space ($\langle r^{2}\rangle$) as a function of time under different periodic drivings for Holstein model. a) We fix the driving frequency $\Omega=0.5$, and change the driving amplitude $J_{1}$ from 00 to $0.8$. We see $\langle r^{2}\rangle$ increases with enhancing the driving amplitude $A$. b) We fix the driving amplitude $J_{1}=0.4$, and change driving frequencies $\Omega$ from $0.2$ to $2$. We see $\langle r^{2}\rangle$ increases first and then decreases to a point where there is no driving. Here, $J_{0}=0.3$, $\omega=0.3$, $g^{2}=0.25$, $T=0.5$, and $N=10$. The distance unit corresponds to $5$, and the time unit corresponds to $82.7$ $fs$.
Figure 5: Mean square displacement in real space ($\langle r^{2}\rangle$) as a function of time under different periodic drivings for Holstein model. a) We fix the driving frequency $\Omega=0.5$, and change the driving amplitude $J_{1}$ from 00 to $0.8$. We see $\langle r^{2}\rangle$ increases with enhancing the driving amplitude $A$. b) We fix the driving amplitude $J_{1}=0.4$, and change the driving frequencies $\Omega$ from $0.2$ to $2$. We see $\langle r^{2}\rangle$ increases first and then decreases to a point where there is no driving. Here, $J_{0}=0.3$, $\omega=0.3$, $g^{2}=0.25$, $T=0.5$, and $N=10$. The distance unit corresponds to $5$, and the time unit corresponds to $82.7$ $fs$.
@@ III.3 Truncation in Brillouin zone
The advantage of using a real-space approach lies in its ability to precisely capture local interactions within a specific spatial region while keeping computational demands in check by limiting the scope of local interactions considered. Similarly, for phenomena restricted to a certain range within the Brillouin zone, a comparable strategy can be employed in reciprocal space. In scenarios where electronic carriers are localized due to either the short time scales not permitting full Brillouin zone traversal or their stabilization at band minima, such confinement can be effectively modeled. This concept has been effectively employed in the theoretical analysis of exciton and trion states in monolayer transition-metal dichalcogenides.Qiu, da Jornada, and Louie (2016) In these studies, the analysis was focused around the K points in the Brillouin zone by implementing a specific truncation radius, thereby streamlining the computations.
In FMD and FSH dynamics, a reciprocal space formulation allows for effective truncation when the initial population is concentrated at or near $k=0$. Here we set $N=20$, $J_{0}=0.6$, $\hbar\omega=0.2$, $J_{1}=0.3$, $\Omega=0.5$, $g^{2}=5$, and $T=1$. We symmetrically truncate the corner of the Brillouin zone and keep $4/5$, $3/5$, and $2/5$ sizes of the original one, and the corresponding size numbers $N$ are $16$, $12$, and $8$, respectively. We emphasize again that truncation in reciprocal space does save a lot of time computational time. Theoretically speaking, the computational time of surface hopping or Ehrenfest dynamics scales as $O(N^{3})$ as a function of $N$, where $N$ is the number of the quantum states. In our simulation, we find that we can use as less as 60% of the total $k$ states in both FSH and FMD methods to recover the results without any truncation. Such that we can save 80% of the total computational time.
@@
Figs. 6 and 7 illustrate the temporal evolution of the whole population (Figs. 6 (a) and 7 (a)) and $P_{0}$ (Figs. 6 (b) and 7 (b)) for the Holstein model under different truncation scales, as calculated by FMD and FSH respectively. Remarkably, these methods yield accurate electronic population profiles even with significant truncation, such as at $k_{0}=\frac{2}{5}\pi$, closely matching the untruncated results, as can be seen in the Figs. 6 (c) and 7 (c).
@@
Figure 6: Truncated Electronic Population Dynamics $P_{k}(t)$ for the Holstein Model Using the FMD Method. a) Comparison of the truncated overall population dynamics over time against the untruncated results. b) Time evolution of the truncated $P_{0}(t)$ at the center of the Brillouin zone, juxtaposed with the untruncated counterpart. c) Analysis of the error $\Delta P_{0}(t)$, quantifying the deviation from the full Brillouin zone results, across various truncation radii $k_{0}$. Parameters: $J=0.6$, $\omega=0.2$, $J_{1}=0.3$, $\Omega=0.5$, $g^{2}=5$, $T=1$, and $N=20$. Here, the time unit corresponds to $82.7$ $fs$.
@@
Figure 7: Truncated Electronic Population Dynamics $P_{k}(t)$ for the Holstein Model Using the FSH Method. a) Comparison of the truncated overall population dynamics over time against the untruncated results. b) Time evolution of the truncated $P_{0}(t)$ at the center of the Brillouin zone, juxtaposed with the untruncated counterpart. c) Analysis of the error $\Delta P_{0}(t)$, quantifying the deviation from the full Brillouin zone results, across various truncation radii $k_{0}$. Parameters: $J_{0}=0.6$, $\omega=0.2$, $J_{1}=0.3$, $\Omega=0.5$, $g^{2}=5$, $T=1$, and $N=20$. Here, the time unit corresponds to $82.7$ $fs$.
@@
However, as observed in Fig. 8, the accuracy of large truncations ($k_{0}=\frac{2}{5}\pi$) in the FMD method for the Peierls model is less satisfactory. In contrast, as shown in Fig. 9, the FSH method demonstrates superior performance under similar truncation conditions, highlighting its advantages. Importantly, truncation calculations is particularly beneficial when dealing with large systems, especially for the Floquet Hamiltonian, making it an attractive strategy for efficiently exploring electronic dynamics in complex models.
@@
Figure 8: Truncated Electronic Population Dynamics $P_{k}(t)$ for the Peierls Model Using the FMD Method. a) Comparison of the truncated overall population dynamics over time against the untruncated results. b) Time evolution of the truncated $P_{0}(t)$ at the center of the Brillouin zone, juxtaposed with the untruncated counterpart. c) Analysis of the error $\Delta P_{0}(t)$, quantifying the deviation from the full Brillouin zone results, across various truncation radii $k_{0}$. Parameters: $J=0.6$, $\omega=0.2$, $J_{1}=0.3$, $\Omega=0.5$, $g^{2}=5$, $T=1$, and $N=20$. Here, the time unit corresponds to $82.7$ $fs$.
@@
Figure 9: Truncated Electronic Population Dynamics $P_{k}(t)$ for the Peierls Model Using the FSH Method. a) Comparison of the truncated overall population dynamics over time against the untruncated results. b) Time evolution of the truncated $P_{0}(t)$ at the center of the Brillouin zone, juxtaposed with the untruncated counterpart. c) Analysis of the error $\Delta P_{0}(t)$, quantifying the deviation from the full Brillouin zone results, across various truncation radii $k_{0}$. Parameters: $J=0.6$, $\omega=0.2$, $J_{1}=0.3$, $\Omega=0.5$, $g^{2}=5$, $T=1$, and $N=20$. Here, the time unit corresponds to $82.7$ $fs$.
@@ IV Conclusions
Our study successfully constructs Floquet Hamiltonian for the Holstein and Peierls models, showing the huge impact of Floquet driving on Floquet replicas. The derived FMD and FSH, explored in both real and reciprocal space, yield consistent results through numerical simulations. This agreement validates the reliability of our theoretical framework, emphasizing the equivalence of real and reciprocal space outcomes. The advantage of simulations in real space is that the algorithm is simple and easy to implement. In addition, simulations in real space are generally more suitable for disordered solids, where the periodicity in real space is broken such that the Block theory is invalid. On the contrary, simulations in reciprocal space provides more physical insights on exciton dynamics in band structure. The band structure can be obtained from electronic structure theory, such that the dynamics in reciprocal space can combine with electronic structure calculation directly. Furthermore, the truncation can be done for simulations in reciprocal space, which saves a lot of computational time. However, FSH is inefficient when dealing with slow drivings due to the large Floquet replicas. The applications of FMD are limited but efficient for large system.